# Visual Autoregressive Modeling for Instruction-Guided Image Editing

### v1

@@ Appendix A Appendix
Training Dataset Processing Details
Our training dataset for VAREdit is constructed from the SEED-Data-Edit (Ge et al. 2024) and ImgEdit (Ye et al. 2025) datasets, yielding a final collection of 3.92 million paired samples. We first extract all single-turn samples of the two datasets and decompose multi-turn conversations in SEED-Data-Edit into single-turn editing pairs for the collection. We identify a significant number of noisy samples within the initial SEED-Data-Edit, characterized by poor instruction alignment, visual blurring or abnormal distortion. To address this, we filter these low-quality samples using a qualified open-source visual language model (VLM), Kimi-VL-A3B-Thinking (Team et al. 2025). The filtering prompt is shown in Figure 7. We discard approximately 1 million samples for which the VLM provided a negative judgment, resulting in our final refined dataset.
@@
Additional Details of Evaluation Metrics
CLIP-based Evaluation
For conventional CLIP-based metrics, we adhere to the established configurations of previous approaches. On the EMU-Edit (Sheynin et al. 2024) benchmark, we follow UltraEdit (Zhao et al. 2024a) to use openai/clip-vit-base-patch32 as the base model. On the PIE-Bench (Ju et al. 2023) benchmark, we use openai/clip-vit-large-patch14 as the base model, which is explicitly specified in the benchmark.
GPT-based Evaluation
Conventional metrics often fail to capture important editing quality aspects, such as spurious modifications or incomplete edits. To address this, we adopt the evaluation protocol from OmniEdit (Wei et al. 2025) to use GPT-4o (Hurst et al. 2024) as an automated judge to provide GPT-Success (Suc.) and GPT-Overedit (Over.) scores. The evaluation prompt is shown in Figure 8. We also introduce GPT-Balance (Bal.) as a balanced evaluation metric for overall editing performance, which stands for the harmonic mean of GPT-Suc. and GPT-Over. metrics. The GPT-Bal. score of an editing sample $\mathbf{e}$ is calculated as:
$\operatorname{GPT-Bal.}(\mathbf{e})=\frac{2\times\operatorname{GPT-Suc.}(\mathbf{e})\times\operatorname{GPT-Over.}(\mathbf{e})}{\operatorname{GPT-Suc.}(\mathbf{e})+\operatorname{GPT-Over.}(\mathbf{e})},$ (8)
and it takes zero value if either the edit is absolutely unsuccessful ($\operatorname{GPT-Suc.}(\mathbf{e})=0$) or severe overediting occurs ($\operatorname{GPT-Over.}(\mathbf{e})=0$). The final reported score for each GPT-based metric on a dataset is the arithmetic mean of the scores across all samples.
@@
Details for Compared Approaches
@@
We assess VAREdit against several tuning-based methods:
@@
  • InstructPix2Pix (Brooks et al. 2023) is the seminal study that introduces the natural, humanized instruction-guided image editing task. It constructs considerable paired editing samples to train an end-to-end editing model based on SD v1.5, which establishes a powerful channel-wise conditional concatenation paradigm in instruction-guided image editing.
@@
  • UltraEdit (Zhao et al. 2024a) introduces a large-scale dataset covering broader instruction types, real images and region editing. Canonical diffusion models trained on the UltraEdit set new benchmarking records. We adopt the published SD v3 variant for comparison.
@@
  • OmniGen (Xiao et al. 2025) is a diffusion-based architecture that tokenizes texts and images into token embeddings and denoises the noise tokens to reconstruct final latent representations through Transformer-based rectified flow optimization. It supports unified image generation, including image editing, subject-driven generation and visual-conditional generation.
@@
  • AnySD (Yu et al. 2025) is another diffusion-based architecture integrating a task-aware routing strategy with mixture-of-experts (MoE) blocks to meet task-specific editing requirements. This model is trained based on SD v1.5 with a newly proposed comprehensive image editing dataset, AnyEdit.
@@
  • EditAR (Mu et al. 2025) is an AR-based architecture including a VQ-autoencoder and a token-by-token autoregressive model to predict target image tokens in latent space. To inject general visual knowledge, EditAR introduces a distillation loss from a vision foundation model feature encoder as a regularization term, thereby improving the text-to-image alignment.
@@
  • ACE++ (Mao et al. 2025) is an instruction-based diffusion framework, which adapts the long-context condition unit module to scenarios of multiple conditions. This model first experiences a zero-reference training stage and then an N-reference training stage to enable support for general instructions across various tasks.
@@
  • ICEdit (Zhang et al. 2025) leverages large-scale Diffusion Transformers to construct an in-context editing framework. It integrates a LoRA-MoE hybrid tuning strategy to enhance flexible adaptation for specialized condition feature processing, while a VLM-based early filtering module is utilized to identify good samples at the very first denoising steps.
@@
All the compared approaches are implemented using the officially released open-source codes with their published default parameter settings.
Additional Implementation Details
@@
Our VAREdit models are built on a foundational VAR-based model, Infinity (Han et al. 2025). We initialize our models with the released pre-trained 2.2B and 8.4B checkpoints for full fine-tuning. For visual tokenization, we use the pre-trained multi-scale tokenizers integrating binary spherical quantization (Zhao et al. 2024b), and the codebook dimension is $32$ and $56$ for the 2.2B and 8.4B models, respectively. During the training stage of $256\times 256$ resolution, all the source and target images will be tokenized into 7 scales with the following list of grid sizes $(h_{k},w_{k})$:
(1,1),(2,2),(4,4),(6,6),(8,8),(12,12),(16,16).(1,1),(2,2),(4,4),(6,6),(8,8),(12,12),(16,16). (9)
@@
As for the training of $512\times 512$ resolution, there will be 10 scales, consisting of the above 7 listed scales and an additional 3 larger shapes: $(20,20),(24,24),(32,32)$. In particular, we only extract the finest-scale tokens of the source images in the finest-scale conditioning strategy. We employ 2D-RoPE to encode the spatial information of all input tokens. For a target token at position $(i,j)$ in scale $k$, its 2D-RoPE coordinate $(x,y)$ is calculated as:
@@
Additionally, we also use the pre-trained scale positional embeddings to mitigate inter-scale confusion.
Numerical Results for Categorical Evaluation
@@
Table 3, 4, 5, 6 present the detailed numerical results of the fine-grained categorical scores shown in radar charts of the main context. It can be observed that VAREdit-2.2B outperforms the compared approaches across the vast majority of editing types, such as addition, removal, coloring and material transfer. VAREdit-8.4B enjoys further enhanced performance in most categories. According to these tables, the 2.2B model shows some limitations in challenging scenarios, including global editing and text editing. Through scaling up the model size and training iterations, the 8.4B model substantially mitigates or even closes these gaps, especially in text editing that achieves 215%+ improvements over the 2.2B variant. These findings strongly demonstrate that our VAREdit is not only effective but also highly scalable. As a result, our VAREdit models exhibit high robustness across various editing types to achieve precise editing results.

### v2

@@ Appendix A Appendix
A.1 Training Dataset Processing Details
Our training dataset for VAREdit-2B is constructed from the SEED-Data-Edit (Ge et al., 2024) and ImgEdit (Ye et al., 2025) datasets, yielding a final collection of 3.92 million paired samples. We first extract all single-turn samples of the two datasets and decompose multi-turn conversations in SEED-Data-Edit into single-turn editing pairs for the collection. We identify a significant number of noisy samples within the initial SEED-Data-Edit, characterized by poor instruction alignment, visual blurring or abnormal distortion. To address this, we filter these low-quality samples using a qualified open-source visual language model (VLM), Kimi-VL-A3B-Thinking (Team et al., 2025). The filtering prompt is shown in Figure 8. We discard approximately 1 million samples for which the VLM provided a negative judgment, resulting in our final refined dataset. For VAREdit-8B, the training dataset includes a large-scale proprietary editing samples.
@@ A.1 Training Dataset Processing Details
A.2 Additional Details of Evaluation Metrics
A.2.1 CLIP-based Evaluation
For conventional CLIP-based metrics, we adhere to the established configurations of previous approaches. On the EMU-Edit (Sheynin et al., 2024) benchmark, we follow UltraEdit (Zhao et al., 2024) to use openai/clip-vit-base-patch32 as the base model. On the PIE-Bench (Ju et al., 2023) benchmark, we use openai/clip-vit-large-patch14 as the base model, which is explicitly specified in the benchmark.
A.2.2 GPT-based Evaluation
Conventional metrics often fail to capture important editing quality aspects, such as spurious modifications or incomplete edits. To address this, we adopt the evaluation protocol from OmniEdit (Wei et al., 2025) to use GPT-4o (Hurst et al., 2024) as an automated judge to provide GPT-Success (Suc.) and GPT-Overedit (Over.) scores. The evaluation prompt is shown in Figure 9. We also introduce GPT-Balance (Bal.) as a balanced evaluation metric for overall editing performance, which stands for the harmonic mean of GPT-Suc. and GPT-Over. metrics. The GPT-Bal. score of an editing sample $\mathbf{e}$ is calculated as: $\operatorname{GPT-Bal.}(\mathbf{e})=\frac{2\times\operatorname{GPT-Suc.}(\mathbf{e})\times\operatorname{GPT-Over.}(\mathbf{e})}{\operatorname{GPT-Suc.}(\mathbf{e})+\operatorname{GPT-Over.}(\mathbf{e})}$, and it takes zero value if either the edit is absolutely unsuccessful ($\operatorname{GPT-Suc.}(\mathbf{e})=0$) or severe overediting occurs ($\operatorname{GPT-Over.}(\mathbf{e})=0$). The final reported score for each GPT-based metric on a dataset is the arithmetic mean of the scores across all samples.
@@ A.2.2 GPT-based Evaluation
A.2.3 Additional Evaluation Details
For ImgEdit-Bench, we follow the official protocol that uses GPT-4o to assign 1-5 ratings, averaged across the dimensions on instruction adherence, image-editing quality, and detail preservation. For GEdit-Bench, we also use the official protocol that utilizes GPT-4.1 to assign 0-10 ratings on semantic consistency (SC), perceptual quality (PQ) and finally calculate the overall score (O) accordingly.
A.3 Details for Compared Approaches
@@
We assess VAREdit against several basic tuning-based methods:
@@
  • InstructPix2Pix (Brooks et al., 2023) is the seminal study that introduces the natural, humanized instruction-guided image editing task. It constructs considerable paired editing samples to train an end-to-end editing model based on SD v1.5, which establishes a powerful channel-wise conditional concatenation paradigm in instruction-guided image editing.
@@
  • UltraEdit (Zhao et al., 2024) introduces a large-scale dataset covering broader instruction types, real images and region editing. Canonical diffusion models trained on the UltraEdit set new benchmarking records. We adopt the published SD v3 variant for comparison.
@@
  • OmniGen (Xiao et al., 2025) is a diffusion-based architecture that tokenizes texts and images into token embeddings and denoises the noise tokens to reconstruct latent representations through Transformer-based rectified flow optimization. It supports unified image generation, including image editing, subject-driven and visual-conditional generation.
@@
  • AnySD (Yu et al., 2025) is another diffusion-based architecture integrating a task-aware routing strategy with mixture-of-experts (MoE) blocks to meet task-specific editing requirements. This model is trained based on SD v1.5 with a newly proposed comprehensive image editing dataset, AnyEdit.
@@
  • EditAR (Mu and others, 2025) is an AR-based architecture including a VQ-autoencoder and a token-by-token autoregressive model to predict target image tokens in latent space. To inject general visual knowledge, EditAR introduces a distillation loss from a vision foundation model as a regularization term, thereby improving the text-to-image alignment.
@@
  • ACE++ (Mao et al., 2025) is an instruction-based diffusion framework, which adapts the long-context condition unit module to scenarios of multiple conditions. This model first experiences a zero-reference training stage and then an N-reference training stage to enable support for general instructions across various tasks.
@@
  • ICEdit (Zhang et al., 2025b) leverages large-scale Diffusion Transformers to construct an in-context editing framework. It integrates a LoRA-MoE hybrid tuning strategy to enhance flexible adaptation for specialized condition feature processing, while a VLM-based early filtering module is utilized to identify good samples at the very first denoising steps.
In addition, we also include four frontier image editing approaches:
  • Step1X-Edit (Liu et al., 2025) is a unified image editing model that combines the robust semantic reasoning of MLLM with a DiT-style diffusion architecture. It maintains a good balance between conditioned image reconstruction and editing instruction following.
  • FLUX.1 Kontext (Labs et al., 2025) is a generative flow matching approach that handles both local editing and generative in-context tasks within a DiT-style architecture.
  • Qwen-Image-Edit (Wu et al., 2025a) is built upon Qwen-Image, simultaneously feeding the input image into Qwen2.5-VL and the VAE Encoder to obtain the representation, serving as conditions of a double-stream MMDiT architecture for effective editing.
  • GPT-4o-Image (Hurst et al., 2024) is a closed-sourced visual generation model that supports generating highly realistic images that accurately follow instructions through simple text descriptions or image uploads.
@@
All the compared approaches (except for GPT-4o-Image) are implemented using the officially released open-source codes with their published default parameter settings.
A.4 Additional Implementation Details
@@
Our VAREdit models are built on a foundational VAR-based model, Infinity (Han et al., 2025). We initialize our models with the released pre-trained 2B and 8B checkpoints for full fine-tuning. For visual tokenization, we use the pre-trained multi-scale tokenizers integrating binary spherical quantization (Zhao and others, 2024), a lookup-free vector quantization method following:
$\mathcal{Q}_{BSQ}(\mathbf{z})=[q(z_{1});q(z_{2});\ldots;q(z_{d})],\quad q(z_{i})=\frac{1}{\sqrt{d}}\cdot\textbf{1}_{[z_{i}\geq 0]}.$ (8)
We set the codebook dimensions $d$ as $32$ and $56$ for the 2B and 8B models, respectively. During the training stages of $256\times 256$ and $512\times 512$ resolution, all the source and target images will be tokenized into 7 scales and 10 scales with the following lists of grid sizes $(h_{k},w_{k})$, respectively:
  • 256px: (1,1),(2,2),(4,4),(6,6),(8,8),(12,12),(16,16).(1,1),(2,2),(4,4),(6,6),(8,8),(12,12),(16,16).
  • 512px: (1,1),(2,2),(4,4),(6,6),(8,8),(12,12),(16,16),(20,20),(24,24),(32,32).(1,1),(2,2),(4,4),(6,6),(8,8),(12,12),(16,16),(20,20),(24,24),(32,32).
@@
In particular, we only extract the finest-scale tokens of the source images in the finest-scale conditioning strategy. We employ 2D-RoPE (Heo et al., 2024) to encode the spatial information of all input tokens before each self-attention module. For a target token at position $(i,j)$ in scale $k$, its 2D-RoPE coordinate $(x,y)$ is calculated as:
@@
Additionally, we also use the pre-trained scale positional embeddings in the Infinity model to mitigate inter-scale confusion.
A.5 Qualitative Study and Analysis
A.5.1 Common-type Image Editing
To further present the superiority of VAREdit, we conduct additional qualitative analysis across diverse editing scenarios, including object-level modifications (addition, replacement, removal), attribute changes (material, text, posture, style, color) and complex compositional edits. As illustrated in Figure 10, VAREdit-8B exhibits satisfactory editing adherence to instructions and high image fidelity on these samples, which serve as commonly seen editing types.
Figure 10: Qualitative editing samples of VAREdit-8B across numerous common editing types.
A.5.2 Imaginative Image Editing
As VAREdit exhibits promising performance in common-type editing, we further explore some sophisticated editing types, including some unseen or imaginative instructions. We have presented the results in Figure 11. Our model achieves relatively successful results on these challenging edits of artistic style transfer (cyberpunk, Zaha’s design style, and Vango’s starry night) and semantic conceptual synthesis (ice flowing waterfall, artistic sound waves, and chess shadows). These results strongly suggest that our model is not merely memorizing training examples but has developed generalizable editing capabilities.
Figure 11: Qualitative editing samples of VAREdit-8B in imaginative image editing.
A.5.3 Reasoning-based Image Editing
To evaluate the capability of VAREdit in reasoning-based image editing He et al. (2025), we present editing results of three representative image samples in ImgEdit-Bench with implicit hypothetical instructions that demand reasoning. As illustrated in Figure 12, by presenting the instruction with implicit “dirty car”, “evaporation” and “cold weather”, VAREdit-8B catches these points and includes the semantics in the edited samples. While some details can be further improved (“blank cup”, “icy lake”), a more balanced dataset curation will better address this gap.
Figure 12: Qualitative editing samples of VAREdit-8B in reasoning-based image editing.
A.5.4 Multi-turn Image Editing
A more challenging scenario is multi-turn image editing, which demands the model to preserve critical details in multiple editing steps. To address this, we present editing results of three representative image samples in ImgEdit-Bench, each with three continual editing instructions. As illustrated in Figure 13, VAREdit-8B performs well and maintains the overall quality after three editing steps.
Figure 13: Qualitative samples of VAREdit-8B in multi-turn image editing.
Figure 14: Additional editing samples of VAREdit-8B in highly-uncertain image editing.
A.5.5 Highly-uncertain Image Editing
In some scenarios, users may present unclear instructions with ambiguity or even conflicts. To investigate this, we conducted a qualitative study on several representative samples. As shown in Figure 14, unclear concepts of “look better” and “add some style” trigger no editing effect, while the underspecified “make the sky louder” is executed by removing the cloud unexpectedly. For conflicting instructions, VAREdit exhibits different patterns. By switching “red” and “blue”, VAREdit tends to ignore “blue” rather than considering both, illustrating a biased focus in some elements. In contrast, the “happy-and-sad” and “curve-and-straight” have been executed at a somewhat balanced level rather than ignoring one of them. In summary, VAREdit demonstrates a degree of robustness in the face of less-than-ideal instructions, and its behavior patterns (converging to priors, ignoring noise, or compromising in conflict) are predictable.
A.5.6 Failure Case Analysis
To comprehensively explore the limitations of VAREdit, we conduct failure case analysis on both 2B and 8B sizes and present some representative editing results. As illustrated in Figure 15, our investigation reveals that while VAREdit-2B struggles with specific categories (e.g., text generation, motion change), VAREdit-8B largely mitigates these issues through scaling. However, even the larger model encounters difficulties in complex compositional tasks that require global modification and preservation of multiple distinct local elements. For example, VAREdit-8B modifies the market into pottery successfully, yet the shapes of some pottery remain abnormal. Moreover, the cross-arm edit poses difficulty to the 8B model by maintaining the general hand fingers. We attribute this limitation to the current density of processing scales and hypothesize that expanding both the image resolution and training data will resolve these fine-grained failures.
Figure 15: Failure editing case results of VAREdit-2B and VAREdit-8B.
A.5.7 User Cross-Validate Study
While we use CLIP and GPT-based scores as the main evaluation metrics, it is unknown how well these values reflect the user preferences. To mitigate this gap, we conducted a blind, randomized human preference study on the PIE-Bench dataset. For each sample in PIE-Bench, we presented participants with the source image, the text instruction, and the three edited images generated by our model (VAREdit-8B) and two strong baselines (Step1X-Edit, FLUX.1 Kontext (dev)). The three generated images were displayed in a randomized order to avoid positional bias. Participants were asked to select the single best image that satisfies two criteria: (1) Instruction following: how accurately the image reflects the textual instruction, and (2) Image quality: the realism of the edit and the absence of artifacts or over-editing. We collected responses from 5 participants (e.g., graduate students in our field with experience in evaluating generative models). The consistency in the results is expected to provide a strong and valuable signal regarding user preference.
Table 4: Cross-validate user study results on PIE-Bench. The best group results are marked in bold.
Method Avg. Count Avg. Rate (%) GPT-Bal.
Step1X-Edit 210.4 30.1 7.351
FLUX.1 Kontext (dev) 237.4 33.9 6.998
VAREdit-8B (Ours) 252.2 36.0 8.106
We report both the average number of wins and the overall preference rate for each model. As illustrated in Table 4, the three methods received a globally balanced number of wins, while VAREdit-8B achieved the highest preference rate (36.0%). This study confirms that VAREdit’s quantitative strengths translate directly into superior perceptual quality for users.
A.6 Quantitative Results on GEdit-Bench
Table 5: Quantitative results of VAREdit and representative approaches on GEdit-Bench (English).
Method Full Intersection
G_SC G_PQ G_O G_SC G_PQ G_O
InstructPix2Pix 3.335 6.210 3.234 3.296 6.189 3.219
AnySD 3.122 5.865 2.919 3.053 5.882 2.854
OmniGen 6.037 5.856 5.154 5.879 5.871 5.005
Step1X-Edit 7.289 6.962 6.618 7.131 6.998 6.444
GPT-4o-Image 7.867 8.097 7.590 7.743 8.133 7.494
VAREdit-2B 5.121 5.301 4.460 4.992 5.220 4.334
VAREdit-8B 7.292 6.581 6.383 7.092 6.626 6.236
We present quantitative results on GEdit-Bench, reporting the SC score, PQ score and the overall score evaluated by GPT-4.1 on VAREdit and some representative approaches. As illustrated in Table 5, GPT-4o-Image scores the highest, while Step1X-Edit and VAREdit-8B are the leading open-source methods. While the absence of training data with Chinese instructions may influence the editing performance of VAREdit, an improved scaling and balance of the training hours and data volume can mitigate this gap.
A.7 Self-Attention Heatmaps Analysis
We provide the full-layer self-attention heatmaps in Figure 16 as a more comprehensive scale dependency analysis. It can be seen that in the first attention layer, the attention scores distribute broadly on almost all source scales, especially the coarser source scales. While in other layers, the attention patterns become highly localized. Specifically, attention scores between source scales and target scales are in strong diagonal structures in deeper Transformer layers, suggesting that attention is primarily confined to tokens in the spatial neighborhood. In this task, the finest accumulated source scale $\mathbf{F}_{K}^{(src)}$ is sufficient to provide reference information for high-quality image editing. Such observations inspire us to design the SAR module to inject scale-matched conditions into the first self-attention layer, while maintaining the finest-scale conditional setting in deeper layers.
Figure 16: Self-attention heatmaps based on the full-scale setting among all transformer layers.
To investigate the attention pattern of the SAR variant, we visualize self-attention heatmaps from some intermediate layers. As illustrated in Figure 17, the first layer exhibits a clear scale-aligned behavior. The attention scores ($\mathbf{Q}_{k}^{(tgt)}\mathbf{K}_{k}^{(ref)\top}/\sqrt{d}$) assigned by each target scale $k$ to the corresponding reference features are predominantly concentrated, without focus on the finest-scale source features. In subsequent layers, the attention pattern becomes highly localized as the same in the full-scale setting. This behavior mirrors the attention patterns observed in the standard full-scale setting, suggesting a shift from global scale-matching to local refinement.
Figure 17: Self-attention heatmaps of PIE-Bench samples based on the SAR-variant across different layers. Note that in the first layer, the focused zone is organized not strictly aligned with the corresponding positions, since $\mathbf{F}_{k}^{(ref)}$ has a spatial size of $(h_{k},w_{k})$ instead of the largest scale $(h_{K},w_{K})$.
A.8 Numerical Results for Categorical Evaluation
@@
Table 6, 7, 8, 9 present the detailed numerical results of the fine-grained categorical scores shown in the radar charts of the main context. It can be observed that VAREdit-2B outperforms the compared approaches across the vast majority of editing types, such as addition, removal, coloring and material transfer. VAREdit-8B enjoys enhanced performance by a substantial margin that outperforms all competitors. According to these tables, the 2B model shows some limitations in challenging scenarios, including global editing and text editing. Through scaling up the model size and training iterations, the 8B model substantially mitigates these gaps, especially in text editing that achieves 150%+ improvements over the 2B variant. These findings strongly demonstrate that our VAREdit is not only effective but also highly scalable. As a result, our VAREdit models exhibit high robustness across various editing types to achieve precise editing results.